### v1

# Perfect supercurrent diode efficiency in chiral nanotube-based Josephson junctions

Abstract
The supercurrent diode effect (SDE) describes superconducting systems where the magnitude of the superconducting-to-normal state switching current differs for positive and negative current bias. Despite the ubiquity of such diode effects in Josephson devices, the fundamental conditions to observe a diode effect in a Josephson junction and achieve perfect diode efficiency remain unclear. In this work, we analyze the supercurrent diode properties of a chiral nanotube-based Josephson junction within a Ginzburg-Landau theory. We find a diode effect and anomalous phase develops across the junction when a magnetic field is applied parallel to the tube despite the absence of spin-orbit interactions in the system. Unexpectedly, the SDE in the junction is independent of the anomalous phase. Alternatively, we determine a non-reciprocal persistent current that is protected by fluxoid quantization can activate SDE, even in the absence of higher-order pair tunneling processes. We show this new type of SDE can lead to, in principle, a perfect diode efficiency, highlighting how persistent currents can be used to engineer high efficiency supercurrent diodes.
With a growing need for low-power fast electronics with low operational temperatures, a rapid development of non-reciprocal superconducting devices has occurred recently 1, 2. Non-reciprocal effects have been intensely investigated in bulk superconductors 3, 4, 5, 6, 7 as well as in Josephson junctions 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19 (JJs)– two superconducting electrodes weakly coupled by a tunneling barrier, such as a normal metal or insulator. In JJs, the Josephson diode effect (JDE) manifests as a difference in magnitude of positive and negative threshold currents where the device switches from a superconducting to a dissipative state, shown schematically in Fig. 1(a). Despite the effect having been known decades ago in devices with geometric inhomogeneities, the subject has received renewed interest based on superconducting diode effects reported in homogeneous devices where non-reciprocity arises from microscopic interactions. In these systems the diode effect, particularly when no magnetic field is applied, has supported the discovery of exotic states, such as time-reversal symmetry broken superconducting states 20, 21. Other zero-field supercurrent diodes include JJs made of iron-based superconductors 22, JJs of twisted bilayer graphene 23, twisted trilayer graphene 24, obstructed atomic insulator JJs 25, strained PbTaSe2 26, and multiferroic JJs 27. Open questions linger about the nature of the SDE in some of these systems and others, which calls for further theoretical modeling to address the fundamental limits of the SDE.
Figure 1: (a) Schematic of voltage-current curve of a supercurrent diode having a negative diode polarity. (b) Schematic of $\mathcal{C}_{2}$-symmetric Fermi surface. (c) Illustration of chiral nanotube on a square lattice. (d) Cartoon of the ChNt-JJs and the helical persistent current inducing the Josephson diode effect.
Theoretical descriptions of the JDE often heuristically focus on symmetry arguments, namely, broken inversion and time-reversal symmetries 28, 29, 30, 1. Assuming these symmetries are broken, one may consider a common minimal expression for the current-phase relationship (CPR) 2: $I_{s}(\phi)=a\sin(\phi)+b\cos(\phi)+c\sin(2\phi)+d\cos(2\phi).$ Here $\phi$ is the phase across the Josephson junction, and $a,b,c$ and $d$ are real-valued constants describing the weights describing phase-coherent Cooper pair tunneling ($a$ and $b$) and pair co-tunneling ($c$ and $d$) supercurrent channels. The parameters $b$ and $d$ are associated with broken time-reversal symmetry (TRS) in the junction. When all four constants are non-zero and treated as independent parameters, then ubiquitous combinations of $(a,b,c,d)$ result in $I_{c+}\neq|I_{c-}|$, where $I_{c+}=\max\left(I_{s}\right)$ and $I_{c-}=\min\left(I_{s}\right),$ and the Josephson diode effect is realized (see Fig. 1(a)) with an efficiency $\eta=\frac{I_{c+}+I_{c-}}{I_{c+}-I_{c-}}$. Using the minimal form of $I_{s}(\phi)$, it is apparently not possible to have a perfectly efficient supercurrent diode where either $I_{c+}$ or $I_{c-}$ are zero. While ideal diode operation has been identified in the ac limit 31, 32, 33, there remains a question in the dc case: is it possible, in principle, to have a perfectly efficient Josephson diode operating in the dc limit? We answer this question in the affirmative by considering chiral nanotube-based Josephson junctions (ChNt-JJs).
In this work, we present a phenomenological Ginzburg-Landau (GL) theory for a ChNt-JJ with a magnetic field applied along the tube. We consider an anisotropic free energy functional that obeys inversion symmetry and choose periodic boundary conditions along non-high symmetry axes to define a chiral nanotube, see Fig. 1(b-c). We first show that in this system, the ChNt-JJ develops an anomalous phase despite the absence of spin-orbit coupling. The link between the anomalous phase and JDE has been investigated in JJs that break inversion symmetry 34, but the existence and possible origin of an anomalous phase in ChNt-JJs has not been identified until now. We also find a diode effect in the absence of pair co-tunneling, due to a non-reciprocal persistent current which is protected by fluxoid quantization in the ChNt-JJ. In this case, a phase-independent persistent current allows us to place an upper bound on the diode efficiency demonstrating how perfect efficiency can be achieved in ChNt-JJs.
We model a superconducting ChNt with higher-order terms in the free energy functional 35:
$F[\psi]-F[0]$ $=\int_{\Omega}\mathbf{dr}_{0}\left(\alpha|\psi|^{2}+\frac{\beta}{2}|\psi|^{4}+\frac{1}{2m_{0}}|\mathbf{p}_{0}\psi|^{2}+\frac{1}{4m_{0}^{2}\zeta_{0}}{\left|{\mathbf{p}}_{0}^{2}\psi\right|}^{2}\right)+\int_{\Omega}\mathbf{dr}_{0}\frac{1}{2m_{1}}\left(|p_{x0}\psi|^{2}-|p_{y0}\psi|^{2}\right)$
$+\int_{\Omega}\mathbf{dr}_{0}\frac{{\left|p_{x0}^{2}\psi\right|}^{2}+{\left|p_{y0}^{2}\psi\right|}^{2}-\frac{1}{2}{\left|\{p_{x0},p_{y0}\}\psi\right|}^{2}}{4m_{1}^{2}\zeta_{1}}+\int_{\Omega}\mathbf{dr}_{0}\frac{{\left|p_{x0}^{2}\psi\right|}^{2}+{\left|p_{y0}^{2}\psi\right|}^{2}-\frac{3}{2}{\left|\{p_{x0},p_{y0}\}\psi\right|}^{2}}{4m_{2}^{2}\zeta_{2}}.$ (1)
Here $m_{0}<m_{1},m_{2}$, $\zeta_{0},\zeta_{1},\zeta_{2}>0$, $\alpha\propto(T-T_{c})$ and $\beta$ are the usual GL coefficients, and $\mathbf{p}=-i\hbar\left(\nabla-i\frac{2e}{\hbar c}\mathbf{A}\right)$ is the momentum operator. The atomic lattice of the system is defined on a 2D sheet $\Omega\in\mathbb{R}^{2}$. Here the kinetic energy contribution to $F$ has inversion symmetry and $\mathcal{C}_{2}$ rotational symmetry when $1/m_{1}\neq 0$ This corresponds to an elliptical Fermi surface, see Fig. 1(b). The circumferential vector $\mathbf{C}=2\pi R(-\sin\theta,\cos\theta),$ where $R$ is the radius of the nanotube, defines the periodic boundary conditions to apply to the nanotube, see Fig. 1(c). To simplify our analysis, we rotate our real space basis $\mathbf{r}=R(\theta)\mathbf{r}_{0}$ where $R(\theta)$ is the 2x2 rotation matrix. The Ginzburg-Landau equations in this case are (see Supplementary Information for details)
$\mathbf{J}=\frac{2e\hbar}{i}\begin{pmatrix}\rho_{1}-\rho_{4}p_{x}^{2}-\rho_{7}p_{y}^{2}-\rho_{8}p_{x}p_{y}&&\rho_{3}-\frac{\rho_{8}}{2}(p_{x}^{2}-p_{y}^{2})-\frac{\rho_{6}}{2}p_{x}p_{y})\\ \rho_{3}-\frac{\rho_{8}}{2}(p_{x}^{2}-p_{y}^{2})-\frac{\rho_{6}}{2}p_{x}p_{y})&&\rho_{2}-\rho_{4}p_{y}^{2}-\rho_{7}p_{x}^{2}+\rho_{8}p_{x}p_{y}\end{pmatrix}\mathbf{j}$ (2)
$\left[\alpha+\beta{\left|\psi\right|}^{2}-\left(\rho_{1}p_{x}^{2}+\rho_{2}p_{y}^{2}+2\rho_{3}p_{x}p_{y}\right)+\rho_{4}\left(p_{x}^{4}+p_{y}^{4}\right)+\left(\rho_{6}+2\rho_{7}\right)p_{x}^{2}p_{y}^{2}+2\rho_{8}\left(\rho_{x}^{3}p_{y}-p_{x}p_{y}^{3}\right)\right]\psi=0,$ (3)
where $\mathbf{j}=\psi^{*}\nabla\psi-\psi\nabla\psi^{*}-i\frac{4e}{\hbar c}\mathbf{A}|\psi|^{2}.$ The coefficients associated with the reduced $\mathcal{C}_{2}$ symmetry are $\rho_{1}=\mu_{1}\cos^{2}\theta+\mu_{2}\sin^{2}\theta,$ $\rho_{2}=\mu_{2}\cos^{2}\theta+\mu_{1}\sin^{2}\theta,$ and $\rho_{3}=(\mu_{2}-\mu_{1})\sin 2\theta$ where $\mu_{1/2}=(m_{1}\pm m_{0})/(2m_{0}m_{1}).$ Higher order kinetic energy terms in the free energy determine the coefficients $\rho_{4}=(\kappa_{1}+\lambda\cos 4\theta)/2$, $\rho_{6}=-2(\kappa_{2}+\lambda\cos 4\theta)$, $\rho_{7}=(\kappa_{1}+2\kappa_{2}-\lambda\cos 4\theta)/2$, and $\rho_{8}=-\lambda\sin 4\theta$ where $\kappa_{1}=\frac{1}{2m_{0}^{2}\zeta_{0}}+\frac{1}{4m_{1}^{2}\zeta_{1}}$, $\kappa_{2}=\frac{1}{4m_{2}^{2}\zeta_{2}}$, and $\lambda=\frac{1}{4m_{1}^{2}\zeta_{1}}+\frac{1}{2m_{2}^{2}\zeta_{2}}$. When $\theta\mod\frac{\pi}{2}\neq 0$, the nanotube is chiral and $\rho_{3}\neq 0$, causing broken mirror symmetry along the nanotube. Here the superfluid stiffness tensor, relating the supercurrent density $\bf J$ to the condensate current $\bf j$, now has a $\bf p$-dependence.
To establish the conditions for the SDE, we consider an external magnetic field along the nanotube $\mathbf{B}=B_{ext}\mathbf{x}$ which breaks time-reversal symmetry. Periodic boundary conditions lead to $p_{y}$ being a good quantum number $p_{y}^{(n)}=\hbar(\pi R^{2}B_{ext}/\Phi_{0}+n)/R$ for some $n\in\mathbb{Z}$ and where $\Phi_{0}$ is the flux quantum. We restrict ourselves to the case of a small diameter nanotube and the lowest sub-band for simplicity, and henceforth set $p_{y}^{(n)}=p_{y}^{(0)}$. Then the order parameter is
$\psi(x,y)$ $=\psi_{x}(x)e^{ip_{y}^{(0)}y/\hbar},$ (4)
for some complex-valued function $\psi_{x}$. From here, to calculate the supercurrent in a Josephson junction, we need to apply appropriate boundary conditions to the solution to Eq. (3) in the junction. The simplest model for a Josephson junction uses rigid boundary conditions 36
$\psi_{x}(x\leq 0)=\psi_{\infty},\quad\psi_{x}(x\geq L)=\psi_{\infty}e^{i\phi},$ (5)
where the normal region defining the junction is $0\leq x\leq L$, $\phi$ is the phase difference across the junction, and $\psi_{\infty}=\sqrt{-\alpha/\beta}.$ A more rigorous treatment of the boundary conditions for the junction can be made and may be important in 2D and 3D geometries 37, but for our quasi-1D system rigid boundary conditions provide sufficient qualitative accuracy. Assuming $T\lesssim T_{c}$ and a short junction $L\ll\xi$ where $\xi$ is the superconducting coherence length, we can solve Eq. (3) by linearizing in the standard way and find
$\psi_{x}(x)$ $\approx\psi_{\infty}L^{-1}e^{iax}\left(L-x+xe^{i(\phi-aL)}\right),$ (6)
where $a=-\frac{\rho_{3}p_{y}^{(0)}}{\hbar\rho_{1}}$. Then the current-phase relationship of the ChNt-JJ is calculated by solving for $J_{x}$ in Eq. (2) which, due to the anisotropic superfluid stiffness, requires both components of $\bf j$:
$\bf{j}$ $=\begin{pmatrix}2i\frac{\psi_{\infty}^{2}}{L}\sin\left(\phi-aL\right)+2ia\psi_{0}^{2}{\left|L-x+x^{i(\phi-aL)}\right|}^{2}\\ \frac{2i\psi_{0}^{2}p_{y}^{(0)}}{\hbar}{\left|L-x+x^{i(\phi-aL)}\right|}^{2}\end{pmatrix}.$ (7)
Writing $j_{0}=\rho_{1}j_{x}+\rho_{3}j_{y}$ and noting $\partial_{x}j_{0}=0=\partial_{y}j_{0}$, we can re-arrange the expression for $J_{x}$ in Eq. (2) to be
$J_{x}=\frac{e\hbar}{i\rho_{1}}\left[2\rho_{1}j_{0}-(\rho_{8}\rho_{1}-2\rho_{3}\rho_{4})p_{x}^{2}j_{y}\right]$
$+$ $\frac{e\hbar}{i\rho_{1}}\left[(\rho_{8}\rho_{1}+2\rho_{3}\rho_{7})p_{y}^{2}+(\rho_{6}\rho_{1}-2\rho_{3}\rho_{8})p_{x}p_{y}\right]j_{y}.$ (8)
From Eq. (7), we see that $p_{y}j_{y}=0$. Then the CPR is given by
$I_{s}(\phi)$ $=I_{c}\left[\sin\tilde{\phi}+\frac{2\hat{\Phi}}{LR}\gamma^{-1}\left(1-\cos\tilde{\phi}\right)\right].$ (9)
Here $I_{c}=\frac{4e\hbar\rho_{1}\psi_{\infty}^{2}A_{\perp}}{L}$, $\tilde{\phi}=\phi-\phi_{0}$, $\phi_{0}=-\frac{\rho_{3}p_{y}^{(0)}L}{\hbar\rho_{1}},$ $\hat{\Phi}=\pi R^{2}B_{ext}/\Phi_{0}$, and
$\gamma$ $=\frac{(\frac{m_{1}}{m_{0}}+\cos 2\theta)^{2}\csc 2\theta}{2m_{1}(\kappa_{1}-\lambda-2\lambda\frac{m_{1}}{m_{0}}\cos 2\theta)}.$ (10)
The CPR in Eq. (9) is invariant under a $2\pi$-phase shift $I_{s}(\phi)=I_{s}(\phi+2\pi)$, reflecting the fact that changing the phase of the order parameter in either of the leads by $2\pi$ does not change the physical state. The critical current $I_{c}$ is distinguished from the conventional expression by the chiral angle dependence of $\rho_{1}=\rho_{1}(\theta)$.
Let’s first consider the simple case where $\mathcal{O}(p^{4})$ terms in Eq. (1) vanish ($\gamma^{-1}=0$) so that $I_{s}=I_{c}\sin\tilde{\phi}$. Here we have an anomalous phase $\phi_{0}$ where the ChNt-JJ has the minimum of its free energy at $\phi_{0}\neq 0$ and $I_{s}(-\phi)\neq-I_{s}(\phi)$ (allowed under broken time-reversal symmetry). In terms of the anisotropy of the nanotube, $\phi_{0}\propto\sin 2\theta/m_{1},$ and in geometry and magnetic field $\phi_{0}\propto B_{ext}A_{n}$ where $A_{n}$ is the surface area of the nanotube in the junction. The latter relationship implies $\phi_{0}$ is linear in the junction length $L$, similar to short Rashba junctions 36. The presence of an anomalous phase here is due to a purely orbital mechanism 38 rather than the more common spin-orbit mechanism 36, 39.
We can also gain some intuition about the anomalous phase $\phi_{0}$ from an analysis of the velocity $v(p_{x})$. In this case, we have $\Delta v_{x}=v_{x}(p_{x})-v_{x}(-p_{x})=4\rho_{3}p_{y}^{(0)}$ so that
$\phi_{0}$ $=-\frac{\Delta v_{x}L}{4\hbar\rho_{1}}.$ (11)
Prior work on a Rashba nanowire JJ with a magnetic field perpendicular to the current flow showed that the anomalous phase due to the spin-orbit interaction 40 is related to a phase shift $\varphi_{0}\propto\Delta v_{F,\sigma}L$ in the Andreev bound state spectrum, where $\Delta v_{F,\sigma}$ is the difference in Fermi velocities of the two spin channels in the junction. Thus, the anomalous phases in Rashba JJs and ChNt-JJs are directly related to broken chiral symmetry in the condensate velocity. This suggests an anomalous phase can arise when the condensate velocity is non-reciprocal, and the JDE develops when the non-reciprocity in the condensate velocity cannot be gauged away in the condensate wavefunction 41.
Figure 2: (a) $\eta$ versus $\theta$ in polar coordinates for $\hat{\Phi}=0.5$ and $L/R=10$. (b) $\eta$ versus $I_{0}/I_{c}$. $\eta$ versus $\theta$ for $\hat{\Phi}=0,~{}0.1,...,~{}0.5$ with (c) $L/R=10$ and (d) $L/R=1$. Here we used $m_{1}/m_{0}=5$, $\kappa_{1}m_{1}/R^{2}=50$, and $\lambda m_{1}/R^{2}=10$.
Returning to the full solution in Eq. (9), we observe the CPR takes an unconventional bipartite form $I_{s}(\phi)=\tilde{I}_{s}(\phi)+I_{0}$ where $I_{0}$ is independent of $\phi$. The phase-independent term $I_{0}$ represents a persistent current in the chiral tube associated with the supercurrent flowing around the tube due to the Little-Parks effect. We have $I_{0}\propto B_{ext}\sin(2\theta)$ so that the current is non-zero only when the tube is chiral and an external magnetic field is applied. While a magnetic field-induced supercurrent around the tube (i.e. $J_{y}\neq 0$) is conventionally expected, it is surprising that this persistent current contributes to the diode effect directly. The picture for the persistent current in this case is depicted in Fig. 1(d), where the current flows along a helical path around the chiral tube. To preserve the requirement that persistent currents form closed loops in equilibrium, the junction will self-tune to the same anomalous phase ($\phi=\phi_{0}$) so that no net supercurrent flows along the $x$-direction.
[…]
The diode efficiency takes a particularly simple form: $\eta=I_{0}/\tilde{I}_{c},$ where $\tilde{I}_{c}$ is the maximum of $\tilde{I}_{s}(\phi)$. This can be solved analytically:
$\eta$ $=\frac{2\hat{\Phi}/(LR)}{\sqrt{(2\hat{\Phi}/(LR))^{2}+\gamma^{2}}},$ (12)
Here we assumed $\gamma^{-1}\neq 0$; otherwise, the diode effect vanishes ($\eta=0$). A representative calculation of $\eta$ as a function of the chiral angle $\theta$ is shown in Fig. 2(a) showing $\eta$ is suppressed at chiral angles $\theta=n\pi/2$ for $n\in\mathbb{Z}$, consistent with Ref. 35. We also observe suppression of $\eta$ at some chiral angles $\theta_{0}$ (e.g. $\sim 0.18\pi$) where $\gamma^{-1}=0$ for $\cos 2\theta_{0}=\frac{m_{0}(\kappa_{1}-\lambda)}{2\lambda m_{1}}$.
We observe that $\eta$ is independent of temperature. This is due to the two quantities $(I_{c+}+I_{c-})$ and $(I_{c+}-I_{c-})$ having the same temperature scaling. We also see that $\mathrm{sgn}(\eta)=\mathrm{sgn}(B_{ext})$ for $10<\hat{\Phi}<$ 1. Both of these features stand in contrast to the superconducting diode effect predicted using the GL theory for a superconducting chiral nanotube 35 (i.e. in the absence of the junction) where the $\eta$ in that case is sensitive to $T$ and generally changes sign for $10<\hat{\Phi}<$ 1. In Ref. 35, the diode effect is calculated using a phenomenological pair breaking Cooper pair momentum along the nanotube and results in a small $\eta$ ($<0.03$). The result in Eq. (12) ignores a depairing momentum in the superconducting leads. This pair breaking mechanism is expected to play some role in actual experiments, suggesting the diode effect may have some weak $T$-dependence in reality.
Evaluating the upper bounds on $\eta$ we see $\eta\rightarrow 1$ as $I_{c}/I_{0}=\gamma LR/2\hat{\Phi}\rightarrow 0,$ see Fig. 2(b). This limit is, in principle, achieved by optimizing the extrinsic contribution $\hat{\Phi}/LR$ to dominate over the intrinsic contribution $\gamma$ determined by the superfluid stiffness of the chiral nanotube (i.e. $m_{1},m_{2},\kappa_{1},\lambda$). Here $p_{y}^{(0)}$ cannot be arbitrarily enhanced by decreasing $R$ since a larger kinetic energy of the condensate eventually leads to a suppression of the superconducting state 35. Then the key parameter for optimizing $\eta$ is a small junction length $L$, as shown in Fig. 2(c-d). Figure 2(c) presents $\eta$ versus $\theta$ for a moderate junction length $L/R=10$. Here $\eta$ achieves maximum values gradually approaching $0.7$ when $\hat{\Phi}$ is increased to $0.5$. However, when $L/R=1$, Fig. 2(d) shows that $\eta$ quickly approaches nearly perfect diode efficiency ($\eta=1$) as $\hat{\Phi}$ increases beyond $0.2$. We quantify the dominance of this extrinsic contribution in terms of a geometric quantity $\eta=\sin\Theta$ where $\Theta=\arctan(I_{0}/I_{c})$, which implies $\eta$ is bounded to values $|\eta|\leq 1$. Thus, in principle it is possible to achieve perfect diode efficiency without non-equilibrium effects 33, 42, 32. Practically speaking, $\eta$ can only approach the ideal limit since $\frac{m_{1}}{m_{0}}+\cos 2\theta>0$ in this chiral nanotube system since we assumed $m_{0}<m_{1}$ (as is typically the case), but there is not a fundamental restriction against $m_{0}=m_{1}$.
In this work, we presented a GL theory for a ChNt-JJ. We derived a purely orbital anomalous phase that develops across the junction when a magnetic field is applied parallel to the tube. The diode effect is suspected to be intimately connected with the anomalous phase 43, but here we find it to be independent of the SDE. We have also shown the origin of the diode effect here is a non-reciprocal persistent current that is protected by fluxoid quantization and can lead to a nearly perfect diode efficiency. While our analysis is not done for a specific material system, the general arguments are applicable to a wide variety of ChNt-JJs. An ideal setting to test SDE in chiral nanotubes is with single-walled nanotubes since effects related to vortices would be suppressed. It has been demonstrated that a thin flake of NbSe2 can induce superconductivity in carbon nanotubes and enable superconducting effects to be probed at high fields 44. This could be a useful platform for future studies of supercurrent diode effects in chiral nanotubes.
Acknowledgements
J.J.C thanks Catalin D. Spataru, Wei Pan and Enrico Rossi for fruitful discussions. The work at Sandia is supported by a LDRD project. Sandia National Laboratories is a multi-mission laboratory managed and operated by National Technology & Engineering Solutions of Sandia, LLC (NTESS), a wholly owned subsidiary of Honeywell International Inc., for the U.S. Department of Energy’s National Nuclear Security Administration (DOE/NNSA) under contract DE-NA0003525. This written work is authored by an employee of NTESS. The employee, not NTESS, owns the right, title and interest in and to the written work and is responsible for its contents. Any subjective views or opinions that might be expressed in the written work do not necessarily represent the views of the U.S. Government. The publisher acknowledges that the U.S. Government retains a non-exclusive, paid-up, irrevocable, world-wide license to publish or reproduce the published form of this written work or allow others to do so, for U.S. Government purposes. The DOE will provide public access to results of federally sponsored research in accordance with the DOE Public Access Plan.
Appendix A Inclusion of pair co-tunneling
In our analysis of the chiral nanotube so far we have ignored the contribution of Cooper pair co-tunneling. We will now show the combination of an anomalous phase and pair co-tunneling is not a sufficient condition for the JDE in a ChNt-JJ. To incorporate co-tunneling effects, we assume an additional contribution to the free energy in the form
[…]
$f_{kin}(\psi)$ $=\frac{1}{2m_{0}}|\mathbf{p}_{0}\psi|^{2}+\frac{1}{2m_{1}}\left(|p_{x0}\psi|^{2}-|p_{y0}\psi|^{2}\right),$ (14)
where $f_{kin}$ is the kinetic part of pair condensate free energy density and for $\Gamma\in\mathbb{R}$ proportional to the pair co-tunneling probability across the JJ. Then the supercurrent density of a short ballistic chiral nanotube-based JJ is given by
$\mathbf{J}$ $=\frac{2e\hbar}{i}\left(1+2\Gamma f_{kin}(\psi)\right)\begin{pmatrix}\rho_{1}&&\rho_{3}\\ \rho_{3}&&\rho_{2}\end{pmatrix}\mathbf{j}.$ (15)
Importantly, we see the form of Eq. (13) creates the same superfluid stiffness anisotropy in the pair and pair co-tunneling channels. The CPR of the small-diameter tube in this case is
$I_{s}(\phi)=$ $\frac{4e\hbar\rho_{1}\psi_{\infty}^{2}A_{\perp}}{L}\sin\left(\phi+\phi_{0}\right)$
$\cdot\left[1+4\hbar^{4}\Gamma\frac{\psi_{\infty}^{2}\rho_{1}}{L^{2}}\left(1-\cos\left(\phi+\phi_{0}\right)\right)\right]$ (16)
$=$ $~{}I_{c}\left(1+2\delta_{\Gamma}\right)\sin\tilde{\phi}-\delta_{\Gamma}\sin 2\tilde{\phi},$ (17)
where $\delta_{\Gamma}=2\hbar^{4}\Gamma\psi_{\infty}^{2}\rho_{1}/L^{2}$. Clearly, there is no diode effect despite an anomalous phase and pair co-tunneling. This is because by propagating the same anisotropy in the pair current channel to the pair co-tunneling channel, the same anomalous phase emerges in both channels suppressing any interference effect between the two channels. If an isotropic pair co-tunneling term had been includes instead of $f_{kin}$, then there may be some interference between the channels resulting in a JDE but the GL equations requires a numerical solution.